OpenStudy (anonymous):

Type the expressions as radicals.? 1.y 3/5 2.y 5/2 3.q 3/2 4.s 6/5

OpenStudy (anonymous):

@Ashleyisakitty

OpenStudy (anonymous):

@phi

OpenStudy (phi):

use the rule that the top number can be the exponent of the base for example, with \[ y^{\frac{3}{5}} \] make this \[ (y^{3})^{\frac{1}{5}} \] now the 2nd rule, replace the number in the bottom part of the fraction with a radical to that power, like this \[\sqrt[5]{y^3}\]